When the price of a commodity is $.10, an individual buys 40 units of it. When the price falls to $.8 what would be elasticity of demand, if demand increase to (i)44 units, (ii) 50 units ?
1
Expert's answer
2020-09-30T13:04:39-0400

"\bf Answers"

"(a) \space \eta_{p} = -0.4" "\space (Inelastic)"


"(b) \space \eta_{p} = -1" "\space (Unitary)"


"\bf Solutions"

We are given:

"P_{0} = \$10, \\ Q_{0} = 40 \space units, \space and \\ P_{1} = \$8"


"PED (\eta_{p}) = \dfrac {\% \space ∆ \space in \space Quantity \space demanded} {\% \space ∆ \space in \space price}"


"\% \space ∆ \space in \space price = \dfrac {(8-10)}{8} × 100\%"


"= \dfrac {-2}{8} × 100\%"


"= -25\%"


"(a) \space Q_{1} = 44 \space units \\"


"\% \space ∆ \space in \space quantity \space demanded"


"= \dfrac {(44 - 40)}{40} × 100\%"


"= \dfrac {4}{40}×100\%"


"= +10\%"


"\eta_{p} = -\dfrac {10\%}{25\%}"


"=\bf -0.4 \space (Inelastic)"


"(b) \space Q_{1} = 50 \space units"


"\% \space ∆ \space in \space quantity \space demanded"


"= \dfrac {(50-40)}{40}×100\%"


"= \dfrac {10}{40}×100\%"


"= +25\%"


"\eta_{p} = -\dfrac {25\%}{25\%}"


"= \bf -1 (Unitary)"